One side of a rectangle is 8cm longer than the other,the perimeter of the rectangle is 72cm.calculate the width of the rectangle. please dat my gmail account is having problem.pls just send me single to indicate that the question have been solved. please.
Answer by Cromlix(4381)

Hi there,
Make one side of rectangle = 'x'
Other side = x + 8
Perimeter = 2 x Length + 2 x Width
72 = 2(x + 8) + 2(x)
72 = 2x + 16 + 2x
Rearrange:
4x = 72 - 16
4x = 56
x = 14 cm
Length = 22 cm
Width = 14 cm.
Hope this helps :-)